# The non-Abelian geometry, topology, and dynamics of a nonreciprocal Su-Schrieffer-Heeger ladder

### v1

@@ II Symmetries of the complex band structure
This $\eta_{k}$-pseudo-Hermiticity of $H_{k}$ guarantees that the eigenvalues are either purely real or appear in complex conjugate pairs, respectively defining the pseudo-Hermitian symmetric phase and the broken phase, according to whether or not the pair of the right and left column eigenvectors $u^{(r/l)}_{k}$ of $H_{k}$ satisfies the relation $\eta_{k}u^{(r)}_{k}=\lambda_{k}u^{(l)}_{k}$ or equivalently $u^{(r){\dagger}}_{k}\eta_{k}u^{(r)}_{k}=\lambda_{k}$ with $\lambda_{k}$ being a nonzero real number. The corresponding pseudo-Hermiticity operation in real space is highly complex, unlike the previously mentioned symmetric operations, which can be referred to in Appendix B.
Compared with the Hamiltonian $\hat{H}$ of the Hermitian two-leg SSH ladder, the dissipation-induced nonreciprocal hopping in $\hat{H}_{\text{HN}}$ loosens some original symmetries, such as the mirror symmetry and the glide symmetry, to the corresponding pseudo ones, and also generalizes the time-reversal symmetry to the field of complex energies.
When $\chi_{+}\neq\chi_{-}$, although the pseudo-mirror, the pseudo-glide, and the $\eta_{k}$-pseudo-Hermitian symmetries are broken, the time-reversal, the chiral, and the inversion symmetries are preserved. Therefore, the eigenenergies in the composite 3D space still belong to the $D_{2h}$ point group and the gap closing can also occur only at the real or imaginary axes of the complex energy plane. The $D_{2h}$ symmetry of the eigenenergies implies that the ladder possesses other hidden pseudo-Hermitian symmetry instead of the $\eta_{k}$-pseudo-Hermitian symmetries in Eq. (9) according to the statement in Appendix A, although we haven’t found the explicit expression of the pseudo-Hermitian operation.
III Topological phase diagram with various geometries of band structure
To explore the non-Abelian topology (i.e., the topology involving a subspace or multiple bands) of this model, given that $\hat{H}_{\text{NH}}$ respects the chiral symmetry, we define a gauge-invariant winding number under PBCs (which is essentially different from the commonly used gauge-dependent ones, see Appendix C):
where
are two $2\times 2$ matrices of the following block off-diagonal Hamiltonian matrix
obtained from the Hamiltonian matrix (8) by changing the basis to $(\hat{a}^{\dagger}_{k\uparrow}~{}\hat{b}^{\dagger}_{k\downarrow}~{}\hat{b}^{\dagger}_{k\uparrow}~{}\hat{a}^{\dagger}_{k\downarrow})$. Define a function $d_{k}\equiv\det h_{k}^{(+)}=(2\kappa-\chi_{+}\chi_{-})+(\kappa^{2}-\delta^{2}+1)\cos k+i(\kappa^{2}-\delta^{2}-1)\sin k$, which generally forms an ellipse in the complex plane by scanning $k$, the polar angle with respect to the ellipse center, from 00 to $2\pi$. Noting that $\det h_{k}^{(-)}=d_{-k}$, Eq. (11) becomes
$w=\frac{1}{2\pi i}\int_{k\in\text{BZ}}\text{d}\ln d_{k},$ (14)
and the topological phase boundaries are determined by the following formula:
$\delta^{2}$ $=$ $(\kappa\pm 1)^{2}\mp\chi_{+}\chi_{-},$
Since $w$ does not change if one simultaneously inverses the both signs of $\kappa$ and $\chi_{+}\chi_{-}$, one can only plot the topological phase diagram for $\kappa\geq 0$ with respect to $\chi_{+}\chi_{-}$, as shown in Fig. 1(b).
Furthermore, one can analytically diagonalize $H_{k}$ to get four energy bands:
@@ III Topological phase diagram with various geometries of band structure
where the two ‘$\pm$’s are uncorrelated, and the two quantities
@@
are both real. Given the nested form of double-square-root in Eq. (17) with real quantities $\alpha_{k}$ and $\beta_{k}$, the geometry of band structures can be classified according to the complexity of bands (i.e., purely real, purely imaginary, and complex) or gap types (i.e., real line, imaginary line, and point). The corresponding conditions are shown in Table 1.
Band Types $(\forall k)$
Purely real $\beta_{k}\geq 0~{}\&~{}\alpha_{k}\geq\sqrt{\beta_{k}}$
 Purely imaginary $\beta_{k}\geq 0~{}\&~{}\alpha_{k}<-\sqrt{\beta_{k}}$
@@
Gap Types $(\forall k)$
Real line       $\big{(}\beta_{k}\geq 0~{}\&~{}\alpha_{k}>\sqrt{\beta_{k}}\big{)}$
or    $\big{(}\beta_{k}<0~{}\&~{}\alpha_{k}\neq 0\big{)}$
Imaginary line  $\big{(}\beta_{k}\geq 0~{}\&~{}\alpha_{k}<-\sqrt{\beta_{k}}\big{)}$ or $\beta_{k}<0$
@@
Table 1: Classification of the geometry of band structures.
For simplicity but without the loss of typical geometries of band structures, in the following we mainly focus on the more symmetric cases of $|\chi_{+}|=|\chi_{-}|\equiv\chi>0$, where we call $\chi_{+}\chi_{-}=\pm\chi^{2}$ the positive and the negative cases, respectively. Thus, the quantities in Eq. (18) are reduced to
@@
where the “$\pm$” in $\beta_{k}^{(\pm)}$ corresponds to the positive and the negative cases, respectively.
From Eq. (17), we also know that any two bands touch in $k$-space must satisfy one of the following two conditions: (I) $\exists k\in$ BZ, $\beta_{k}=0$, which indicates an always-existing gapless point $k=\pi\,(0)$ for the positive (negative) case, and an inequality
@@
where the ‘$\pm$’ corresponds to the positive and the negative cases, respectively; (II) $\exists k\in$ BZ, $\beta_{k}=\alpha_{k}^{2}$, which leads to the gapless equations for the positive case (i.e., $\chi_{+}\chi_{-}=\chi^{2}$):
$\delta^{2}=\left\{\begin{aligned} &(\kappa-1)^{2}+\chi^{2}\\ &(\kappa+1)^{2}-\chi^{2},~{}~{}\chi^{2}\leq(\kappa+1)^{2}\\ &\,\kappa^{2}-1,~{}~{}\kappa\geq 1~{}\&~{}2(\kappa-1)\leq\chi^{2}\leq 2(\kappa+1),\end{aligned}\right.$ (21)
and for the negative case (i.e., $\chi_{+}\chi_{-}=-\chi^{2}$):
$\delta^{2}=\left\{\begin{aligned} &(\kappa+1)^{2}+\chi^{2}\\ &(\kappa-1)^{2}-\chi^{2},~{}~{}\chi^{2}\leq(\kappa-1)^{2}\\ &\,\kappa^{2}-1,~{}~{}\kappa\leq-1~{}\&~{}2(-\kappa-1)\leq\chi^{2}\leq 2(-\kappa+1).\end{aligned}\right.$ (22)
@@
Clearly, the gapless equations (21) and (22) are just Eq. (16) for the topological phase boundaries. Moreover, along with the up/lower bounds in Eq. (20), the phase diagram can be divided into regions with different geometries of band structures, as shown in Figs. 1(c)-1(r), which can also be derived from the conditions in Table 1; in this model, there is no purely imaginary band structure for the positive and the negative cases. Notably, the lower bound of the inequality (20) is just the boundary of the purely real band structure, which only exists for
$\delta^{2}\leq\frac{\chi^{2}(\kappa\pm 1)^{2}}{\chi^{2}+4},$ (23)
@@
where one can verify that $u^{(r){\dagger}}_{k}\eta_{k}u^{(r)}_{k}$ is a nonzero real number for each $k$ for the positive case, indicating the pseudo-Hermitian symmetric phases for all $k$’s. The symmetric phase in the negative case cannot be straightforwardly verified due to the lack of a simple expression of the pseudo-Hermitian operator, but one can also verify this using the more complex expression in Appendix A.
IV Non-Abelian dynamics under an external constant force in the pseudo-Hermitian symmetric regime
@@
Figure 2: (a) A typical purely real band structure with parameters $(\kappa,\chi^{2},\delta^{2})=(2.5,0.06,0.06)$ in the pseudo-Hermitian symmetric phase of the positive case. The characteristic quantities are $(w_{2},\Delta,w_{4})\approx(0.49,2.97,7.47)$. (b-g) The evolution of unnormalized populations $|\alpha_{n}(t)|^{2}$ in $n$th band given an initial state located at $k=0$ of the lowest band in (a) for various strengths of the force $F=0.05$, $0.5$, $3$, $5$, $10$, and $50$, respectively, which are all governed by Eq. (26). (h,i) The evolution of populations in each band given the same initial state as in (b-g) but using the four-band and two-band Wilson lines $W_{0\rightarrow k}$ and $W_{0\rightarrow k}^{(-)}$, governed by Eqs. (28) and (30), respectively. The numerical calculation is done by dividing the path $0\rightarrow 2\pi$ into $49$ equal pieces, and the details of method can be found in Appendix D.
@@
Since there always exists a band touching point at $k=\pi\,(0)$ for the positive (negative) case, the adiabatic evolution for a single band cannot work any more in $k$-space. Here we focus on the non-Abelian dynamics (i.e., the dynamics involving multiple bands) in the pseudo-Hermitian symmetric phase where all bands are real and thus the dynamics is stable without amplitudes decaying or amplifying in the long-time limit, while in the cases with complex bands, the eigenstates with eigenenergies of the largest imaginary parts will dominate in the long-time limit.
To investigate the non-Abelian dynamics in $k$-space, we add an external constant force $F$ along, say, the right direction of the ladder, and the evolution satisfies the time-independent Schrödinger equation ($\hbar=1$) [7]:
$i{\partial_{t}}|\psi(t)\rangle=(\hat{H}_{\text{HN}}-F\hat{X})|\psi(t)\rangle,$ (24)
@@
where $\partial_{t}\equiv\partial/\partial t$ and for simplicity the position operator $\hat{X}$ is defined only with respect to the positions of unit cells:
$\hat{X}:\hat{a}_{j\sigma}^{({\dagger})}\rightarrow j\hat{a}_{j\sigma}^{({\dagger})},\,\hat{b}_{j\sigma}^{({\dagger})}\rightarrow j\hat{b}_{j\sigma}^{({\dagger})}.$ (25)
@@
The initial state is set as the superposition of Bloch states at $k_{0}$, i.e., $|\psi(0)\rangle=\sum_{n=1}^{4}\alpha_{n}(0)|u_{k_{0}n}^{(r)}\rangle$, where $\alpha_{n}(t)$ is the amplitude of the Bloch right state $|u_{k_{0}n}^{(r)}\rangle$ of $\hat{H}_{\text{NH}}$ in the $n$th band at time $t$, and we normalize the coefficients at $t=0$ by $\sum_{n}|\alpha_{n}(0)|^{2}=1$. With the ansatz that the initial state uniformly evolves to another superposition state $|\psi(t)\rangle=\sum_{n=1}^{4}\alpha_{n}(t)|u_{k(t),n}^{(r)}\rangle$ at $k(t)=k_{0}+Ft$, the solution to Eq. (24) is as follows:
[…]
where $\alpha(t)=(\alpha_{1}(t),\alpha_{2}(t),\alpha_{3}(t),\alpha_{4}(t))^{T}$, $\mathcal{P}$ is the path-ordered operator, and

### v2

@@ II.2 Hamiltonian in $k$ space and the symmetries
but it cannot be easily cast in a simple form for $\chi_{+}\neq\chi_{-}$. This $\eta_{k}$-pseudo-Hermiticity of $H_{k}$ guarantees that the eigenvalues are either real or appear in complex-conjugate pairs, respectively defining the pseudo-Hermitian symmetric phase and the broken phase, according to whether or not the pair of the right- and left-column eigenvectors $u^{(r/l)}_{k}$ of $H_{k}$ satisfies the relation $\eta_{k}u^{(r)}_{k}=\lambda_{k}u^{(l)}_{k}$ or equivalently $u^{(r){\dagger}}_{k}\eta_{k}u^{(r)}_{k}=\lambda_{k}$ with $\lambda_{k}$ being a nonzero real number. It is worth noting that even for $\chi_{+}=\chi_{-}$, where $\eta_{k}$ is explicitly written in a simple form in Eq. (10), the corresponding pseudo-Hermiticity operation in real space is highly complex, which can be referred to in Appendix B.
Based on these underlying symmetries, we analyze the following non-Abelian properties, including the geometry of band structure with $D_{2h}$ symmetry, the topology under the chiral symmetry, and the dynamics in the pseudo-Hermitian symmetric regime.
III Geometry of band structure
The geometry of band structure can be obtained by analytically diagonalizing $H_{k}$, yielding four energy bands:
@@ III Geometry of band structure
where two $\pm$”s are uncorrelated, and the two quantities
@@
are both real. Given the nested form of double-square-root in Eq. (11) with real $\alpha_{k}$ and $\beta_{k}$, the geometry of band structure can be classified by its complex value (i.e., real, purely imaginary, and complex bands) and gap type (i.e., real line, imaginary line, and point gaps). The corresponding conditions are shown in Table 1.
Table 1: Classification of the geometry of band structure.
Complex value $(\forall~k)$
Real $\beta_{k}\geq 0~\text{and}~\alpha_{k}\geq\sqrt{\beta_{k}}$
Purely imaginary $\beta_{k}\geq 0~\text{and}~\alpha_{k}<-\sqrt{\beta_{k}}$
@@
Gap type $(\forall~k)$
Real line $(\beta_{k}\geq 0~\text{and}~\alpha_{k}>\sqrt{\beta_{k}})$
or $(\beta_{k}<0~\text{and}~\alpha_{k}\neq 0)$
Imaginary line  $(\beta_{k}\geq 0~\text{and}~\alpha_{k}<-\sqrt{\beta_{k}})$ or $\beta_{k}<0$
@@
For simplicity but without loss of the typical geometry of band structure, we focus on two symmetric cases of $|\chi_{+}|=|\chi_{-}|\equiv\chi>0$, where we call $\chi_{+}\chi_{-}=\pm\chi^{2}$ the positive and the negative cases, respectively. Typical band structures are shown in Figs. 1(c)–1(r). Thus, the quantities in Eq. (12) are reduced to
@@
where the $\pm$” in $\beta_{k}^{(\pm)}$ corresponds to the positive and the negative cases, respectively.
Equation (11) indicates that the touching of any two bands in $k$ space must satisfy one of the following two conditions: (I) $\exists~k\in$ BZ, $\beta_{k}=0$, which indicates an always-existing gapless point $k=\pi\,(0)$ for the positive (negative) case, and an inequality
@@
where the $\pm$’ corresponds to the positive and the negative cases, respectively; (II) $\exists~k\in$ BZ, $\beta_{k}=\alpha_{k}^{2}$, which leads to the gapless equations:
$\delta^{2}=\left\{\begin{aligned} &(\kappa-1)^{2}+\chi^{2},\\ &(\kappa+1)^{2}-\chi^{2},~\chi^{2}\leq(\kappa+1)^{2},\\ &\,\kappa^{2}-1,~\kappa\geq 1~\text{and}~2(\kappa-1)\leq\chi^{2}\leq 2(\kappa+1)\end{aligned}\right$. (15)
for the positive case (i.e., $\chi_{+}\chi_{-}=\chi^{2}$), and
$\delta^{2}=\left\{\begin{aligned} &(\kappa+1)^{2}+\chi^{2},\\ &(\kappa-1)^{2}-\chi^{2},~\chi^{2}\leq(\kappa-1)^{2},\\ &\,\kappa^{2}-1,~\kappa\leq-1~\text{and}~2(-\kappa-1)\leq\chi^{2}\leq 2(-\kappa+1)\end{aligned}\right$. (16)
@@
for the negative case (i.e., $\chi_{+}\chi_{-}=-\chi^{2}$). These gapless equations along with the upper and the lower bounds in Eq. (14) can divide a phase diagram into different geometric regimes of band structure, as shown in Fig. 1(b), which can alternatively be derived from the conditions in Table 1 (no purely imaginary band structure for the positive and the negative cases in this model). Notably, the lower bound of the inequality (14) is just the boundary of the region of real band structures (i.e., the pseudo-Hermitian symmetric phase), which satisfies
$\delta^{2}\leq\frac{\chi^{2}(\kappa\pm 1)^{2}}{\chi^{2}+4}.$ (17)
@@
In this region, one can numerically verify for the positive case the condition that $u^{(r){\dagger}}_{k}\eta_{k}u^{(r)}_{k}$ is a nonzero real number for each $k$. For the negative case, although a simple expression of $\eta_{k}$ is lacking, one may verify this condition by resorting to the more complex expression of $\eta_{k}$ in Appendix A.
IV Non-Abelian topology under the chiral symmetry
IV.1 Winding number under PBCs
To explore the non-Abelian topology (i.e., the topology involving a subspace of multiple bands), given that $\hat{H}_{\text{NH}}$ respects the chiral symmetry, we define a gauge-invariant winding number under PBCs (which is essentially different from the commonly used gauge-dependent ones, see Appendix C for detailed comparison):
where
are defined in the following off-diagonal block Hamiltonian matrix:
obtained from the Hamiltonian matrix (8) by changing the basis to $(\hat{a}^{\dagger}_{k\uparrow}~\hat{b}^{\dagger}_{k\downarrow}~\hat{b}^{\dagger}_{k\uparrow}~\hat{a}^{\dagger}_{k\downarrow})$. Define a function $d_{k}\equiv\det h_{k}^{(+)}=(2\kappa-\chi_{+}\chi_{-})+(\kappa^{2}-\delta^{2}+1)\cos k+i(\kappa^{2}-\delta^{2}-1)\sin k$, which generally forms an ellipse in the complex plane by scanning $k$, the polar angle with respect to the ellipse center, from 0 to $2\pi$. Noting that $\det h_{k}^{(-)}=d_{-k}$, Eq. (18) becomes
$w=\frac{1}{2\pi i}\int_{k\in\text{BZ}}\text{d}\ln d_{k},$ (21)
and the topological phase boundaries are determined by
$\delta^{2}$ $=$ $(\kappa\pm 1)^{2}\mp\chi_{+}\chi_{-},$
which, for the positive and the negative cases, are identical to the gapless equations (15) and (16), indicating that the topological phase transitions must be accompanied by the gap closing. Since $w$ does not change if one simultaneously inverses both signs of $\kappa$ and $\chi_{+}\chi_{-}$, one can only plot the topological phase diagram for $\kappa\geq 0$ with respect to $\chi_{+}\chi_{-}$, as an example of Fig. 1(b).
IV.2 BBC and critical NHSE
Figure 2: (a) Comparison between the winding number $w$ (black solid line) under PBCs and the open-bulk winding number $w_{r}$ (red crosses) under OBCs along the orange dotted line at $\chi=2$ in Fig. 1(b), where the parameters for calculating $w_{r}$ are $(L,L^{\prime},l)=(200,160,20)$. (b) Energy spectra of $\hat{H}_{\text{NH}}$ in the complex energy plane under OBCs for $L=5,6,\dots,100$ (dots with colors from dark blue to light green) at weak inter-leg coupling $\chi=0.01$ of the positive case. The spectrum (two red solid lines with dots in-between) under OBCs without inter-leg coupling for $L=100$ and the spectrum (gray dashed circle) under PBCs at $\chi=0.01$ of the positive case are also plotted for reference. Insets: Typical density distributions in leg $\uparrow$ (purple) and leg $\downarrow$ (yellow) for the eigenstates indicated by the arrows. The other parameters are $(\kappa,\delta)=(2.5,2)$. (c) The averaged density distributions versus the scaled unit-cell index for $L=5$ (circles), $50$ (triangles), and $100$ (squares) in leg $\uparrow$ (purple) and leg $\downarrow$ (yellow).
To investigate whether the nonreciprocal SSH ladder satisfies the BBC, we numerically plot the spectra under OBCs with $L=400$ unit cells [top layers of Figs. 1(c)–1(r)], and find that except for the topological zero-mode end states, which are protected by the chiral symmetry, the bulk spectra are identical to those under PBCs [bottom layers in Figs. 1(c)–1(r)], indicating the survival of BBC in this ladder. In addition, we also numerically calculate the open-bulk winding number in real space under OBCs [59, 50]:
where $\hat{C}$: $\hat{a}_{j\uparrow}^{({\dagger})}\rightarrow\hat{a}_{j\uparrow}^{({\dagger})}$, $\hat{a}_{j\downarrow}^{({\dagger})}\rightarrow-\hat{a}_{j\downarrow}^{({\dagger})}$, $\hat{b}_{j\uparrow}^{({\dagger})}\rightarrow-\hat{b}_{j\uparrow}^{({\dagger})}$, $\hat{b}_{j\downarrow}^{({\dagger})}\rightarrow\hat{b}_{j\downarrow}^{({\dagger})}$ is the chiral operator, $\hat{Q}\equiv\sum_{n}\big{[}|\phi_{n}^{(r)}\rangle\langle\phi_{n}^{(l)}|-\hat{C}|\phi_{n}^{(r)}\rangle\langle\phi_{n}^{(l)}|\hat{C}\big{]}$ is the “$Q$-matrix” operator constructed by the $n$th chiral-operation paired right/left eigenstates $\big{\{}|\phi_{n}^{(r/l)}\rangle,\hat{C}|\phi_{n}^{(r/l)}\rangle\big{\}}$ of $\hat{H}_{\text{NH}}$ under OBCs, and the primed trace means that the trace is done over a central $L^{\prime}=L-2l$ unit cells out of the whole $L$ unit cells of the ladder with $l$ unit cells being cut off from each end of the ladder to avoid the boundary effects when the numerical calculation is executed. It is shown in Fig. 2(a) that $w_{r}$ is close to $w$ for large systems, which hints that there is no BBC breaking and thus no skin effect in the thermodynamic limit $L\rightarrow\infty$. This can also be reflected by the winding number of energy under PBCs [22, 60]:
However, when the inter-leg coupling vanishes (i.e., $\chi=0$), the ladder is decoupled into two independent nonreciprocal SSH chains, which are well known for the breakdown of BBC and the emergence of NHSEs under OBCs. This indicates that in the thermodynamic limit, the BBC is abruptly restored with the introduction of an infinitesimal inter-leg coupling, which is called the critical NHSE [42]. The critical NHSE leads to a noticeable alteration of the energy spectra and eigenstates under OBCs in the weak inter-leg coupling limit as the system size increases, as shown in Fig. 2(b). When the size is small, the energy spectra and eigenstates are close to those of the decoupled nonreciprocal SSH chains under OBCs, i.e., the spectra are linelike and the eigenstates demonstrate NHSEs, indicating that the coupling effects are negligible; when the size becomes larger, they converge to those under PBCs, i.e., the linelike spectra expand to loops and the eigenstates tend to be delocalized without NHSEs as Bloch waves, indicating that the weak coupling becomes stronger in the thermodynamic limit. Quantitatively, for a finite-size, weakly inter-leg coupled ladder under OBCs, the distribution of bulk eigenstates remains the exponential form $|\phi_{j}|^{2}\propto e^{-|j-j_{b}|/\xi}$ of the skin bulk states of the nonreciprocal SSH chain, localized in the boundary unit cell $j_{b}$, but the decaying length $\xi$ of the amplitude $\phi_{j}$ is size dependent and satisfies the scale-free power law [42, 43]: $\xi(L)\propto L^{c}$, unlike the nonreciprocal SSH chain, where $\xi^{-1}=2^{-1}\ln|\kappa_{+}/\kappa_{-}|$ is size independent [18]. To show the power law, we define the averaged density distributions in leg $\sigma$ ($\sigma=\uparrow$ and $\downarrow$) for certain system size with unit-cell number $L$:
where the amplitudes $a_{j\sigma n}$ and $b_{j\sigma n}$ are defined in the eigenvalue equation $H\phi_{n}^{(r)}=\epsilon_{n}\phi_{n}^{(r)}$ with the $n$th right-column eigenvector $\phi_{n}^{(r)}=(\{a_{j\uparrow n},a_{j\downarrow n},b_{j\uparrow n},b_{j\downarrow n}\})^{T}$. Figure 2(c) shows the averaged density distributions $(\bar{\rho}_{j\uparrow}/\bar{\rho}_{1\uparrow},\bar{\rho}_{j\downarrow}/\bar{\rho}_{L\downarrow})$ normalized by the corresponding localized boundary values versus the scaled unit-cell index $j/L$ for different system sizes. The collapse into the same lines for, respectively, upper and lower legs verifies the scale-free power law with the power $c=1$.
In addition, the critical NHSEs also have impact on the topological end states. Since the zero-mode end state protected by the chiral symmetry appears under OBCs when $\delta^{2}<$(>)>) $\kappa^{2}-1$ for the nonreciprocal SSH chain with the (non)reciprocal unit cells being the boundary cells [18], in the decoupled case the two relatively shifted nonreciprocal SSH chains as a whole always possess end states under OBCs. This means that a topological end state [e.g., the bottom right inset of Fig. 2(b)] in small-sized ladder with weak inter-leg couplings under OBCs may evolve into a bulk state [e.g., the top right inset of Fig. 2(b)] of a topologically trivial phase in the thermodynamic limit, of which, according to the phase diagram in Fig. 1(b), the winding number $w=0$, i.e., the BBC is restored in the thermodynamic limit. At this moment, one may realize that the red vertical line at $\chi^{2}=0$ of the phase diagram in Fig. 1(b) is singular in the thermodynamic limit.
It is apparent that if we add a leg-dependent potential, the NHSE under OBCs along with the breakdown of BBC will return in the thermodynamic limit, with NHSEs depending on the eigenenergies, and thus the so-called bipolar NHSE naturally appears [59, 61].
V Non-Abelian dynamics in the pseudo-Hermitian symmetric regime
@@
Figure 3: (a) A typical real band structure with parameters $(\kappa,\chi^{2},\delta^{2})=(2.5,0.06,0.06)$ in the pseudo-Hermitian symmetric phase of the positive case. The characteristic quantities are $(w_{2},\Delta,w_{4})\approx(0.49,2.97,7.47)$. (b)–(g) The evolution of $|\alpha_{n}(t)|^{2}$ in $k$ space with an initial state being located in the lowest band at $k=0$ in (a) for various strengths of the force $F=0.05$, $0.5$, 33, 55, $10$, and $50$, calculated by Eq. (29). (h),(i) The evolution of $|\alpha_{n}(t)|^{2}$ in $k$ space with the same initial state as in (b)–(g), but calculated by the four-band and two-band Wilson lines $W_{0\rightarrow k}$ and $W_{0\rightarrow k}^{(-)}$ in Eqs. (31) and (33), respectively. All the evolutions are numerically calculated by dividing the path $k:0\rightarrow 2\pi$ into $49$ equal pieces, and the details of the method can be found in Appendix D.
@@
Since there always exists a band-touching point at $k=\pi\,(0)$ for the positive (negative) case, the adiabatic evolution for a single band cannot work any more in $k$ space. Here we focus on the non-Abelian dynamics (i.e., the dynamics involving multiple bands) in the pseudo-Hermitian symmetric regime where all bands are real and thus the dynamics is expected stable without amplitudes decaying or amplifying in the long-time limit, while in the cases with complex bands, the eigenstates with eigenenergies possessing the largest imaginary parts will dominate in the long-time limit.
@@
To investigate the non-Abelian dynamics in $k$ space, we add an external constant force $F$ along the ladder, i.e., the Hamiltonian in Eq. (1) becomes $\hat{H}-F\hat{X}$, where for simplicity the position operator $\hat{X}$ is defined only with respect to the positions of unit cells, i.e.,
$\hat{X}=\sum_{j\sigma}j(\hat{a}_{j\sigma}^{{\dagger}}\hat{a}_{j\sigma}+\hat{b}_{j\sigma}^{{\dagger}}\hat{b}_{j\sigma}).$ (27)
Following the similar derivation from Eq. (4) to (5), the short-time dynamics before a quantum jump occurs can be well captured by the time-dependent non-Hermitian Schrödinger equation ($\hbar=1$):
$i{\partial_{t}}|\psi(t)\rangle=(\hat{H}_{\text{NH}}-F\hat{X})|\psi(t)\rangle,$ (28)
@@
where $\partial_{t}\equiv\partial/\partial t$, and the loss term $i\delta\hat{N}$ in Eq. (5), which only contributes an overall decaying factor to the dynamics, is ignored. The initial state is set as a superposition of Bloch states of different bands at $k_{0}$, and presumably evolves to another superposition at $k(t)=k_{0}+Ft$ at time $t$ [7], i.e., $|\psi(t)\rangle\approx\sum_{n=1}^{4}\alpha_{n}(t)|u_{k(t),n}^{(r)}\rangle$, where $\alpha_{n}(t)$ is the amplitude of right Bloch state $|u_{k(t)n}^{(r)}\rangle$ of $\hat{H}_{\text{NH}}$ in the $n$th band at $k(t)$; we normalize the amplitude at $t=0$, i.e., $\sum_{n}|\alpha_{n}(0)|^{2}=1$. Thus, the solution to Eq. (28) reads as
$\alpha(t)=\mathcal{P}\exp\Big{\{}-i\int_{k_{0}}^{k}dk\big{(}\Lambda_{k}/F+A_{k}\big{)}\Big{\}}\alpha(0),$ (29)
where $\alpha(t)=[\alpha_{1}(t),\alpha_{2}(t),\alpha_{3}(t),\alpha_{4}(t)]^{T}$, $\mathcal{P}$ is the path-ordered operator, and